OpenStudy (anonymous):

Can someone please help me? Verify each trigonometric equation by substituting identities to match the right hand side of the equation to the left hand side of the equation. cot x sec4x = cot x + 2 tan x + tan3x

OpenStudy (anonymous):

This is as far as ive gotten\[\frac{ 1 }{ sinxcos^3x}\]

OpenStudy (anonymous):

I dont understand how he got cos x on top

OpenStudy (anonymous):

can someone help me

OpenStudy (whpalmer4):

Which one are you working on?

OpenStudy (anonymous):

For the third one: 1 + sec2x sin2x = sec2x Remember:\[\sec(2x)=\frac{ 1 }{ \cos(2x) }\]Knowing this:\[\frac{ 1 }{ \cos(2x) }*\sin(2x)=\frac{ \sin(2x) }{ \cos(2x) }=\tan(2x)\]So we have now:\[1+\tan(2x)=\sec(2x)\]Which is a general identity.

OpenStudy (loser66):

\[sec^2x or sec(2x)\]

OpenStudy (anonymous):

I actually need help with the first 1

OpenStudy (anonymous):

For the fifth one: - tan2x + sec2x = 1 Just move the tan2x over to the right side, then we'll get a general identity (Just like the one I posted above.) \[1+\tan(2x)=\sec(2x)\]

OpenStudy (loser66):

again, \[sec^4x\]or sec(4x)?

OpenStudy (anonymous):

i didnt notice i put all the problems up there

OpenStudy (anonymous):

cot x sec4x = cot x + 2 tan x + tan3x

OpenStudy (loser66):

hey, you didn't know, how can others help? they are different, completely. pal. You must post the right problems, please.

OpenStudy (anonymous):

Alright sorry